If the confidence interval is given as 90%, 93%, 96% and 98%. Find the value of z. Round the answer to the two decimal places.


1
Expert's answer
2022-05-04T14:30:33-0400

The critical value for "\alpha = 0.1" is "z_c = z_{1-\alpha/2} = 1.64."

The critical value for "\alpha = 0.07" is "z_c = z_{1-\alpha/2} = 1.81."

The critical value for "\alpha = 0.04" is "z_c = z_{1-\alpha/2} = 2.05."

The critical value for "\alpha = 0.02" is "z_c = z_{1-\alpha/2} = 2.33."